OpenStudy (anonymous):

Solve each equation. Write all proposed solutions. Cross out those that are extraneous.

OpenStudy (anonymous):

\[\sqrt{b^2+b} = \sqrt{3-b^2}\]

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

b^2+b^2=3^2-b^2 b^2=9 b=sqrt(9) b=3